OpenStudy (anonymous):

what is 3over 5 times 4 over 7 explain

OpenStudy (anonymous):

I was explaining. You wandered over here.

OpenStudy (anonymous):

It's the same thing you were doing on the last problem.

OpenStudy (anonymous):

Turn the one on the bottom upside down and multiply it instead (Polpak has another method).

OpenStudy (anonymous):

Turning fractions = the devil (imo)

OpenStudy (anonymous):

Do u see what I mean?

OpenStudy (anonymous):

Actually, I just noticed that we have both read the question wrong,lol.

OpenStudy (anonymous):

3/ 5 * 4/7 is straightforward, multiply the tops and the bottoms. To divide the second into the first, invert the second and multiply. So 3/5 * 7/4

OpenStudy (anonymous):

\[\frac{a}{b}\times\frac{c}{d} = \frac{a\times c}{b\times d}\]

OpenStudy (anonymous):

12over 35 right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

wuts answer

OpenStudy (anonymous):

Your answer is right.